OpenStudy (anonymous):

Please help! The suspension cables between the two towers of the Mackinac Bridge in Michigan form a parabola that can be modeled by the graph of y = 0.000097x2 − 0.37x + 540 where x and y are measured in feet. What is the height of the cable above the water at its lowest point?

OpenStudy (anonymous):

@whpalmer4 can you please help me

OpenStudy (whpalmer4):

Here's the formula for finding the vertex of a parabola in the form \[y = ax^2+bx+c\] \[x = -\frac{b}{2a}\]\[y = a(\frac{b^2}{4a^2}) + b(\frac{-b}{2a}) + c = \frac{b^2}{4a}-\frac{b^2}{2a}+c = -\frac{b^2}{4a}+c\]or just plug \(x\) into the original formula

OpenStudy (anonymous):

@whpalmer4 thank you!